OpenStudy (anonymous):

Simplest form of (x^3/x^-1)^1/4

OpenStudy (anonymous):

steps please. i know the answers

OpenStudy (lgbasallote):

simplify the one in the parenthesis first. that's[(x^3)(x^1)] ^1/4 so [x^4]^1/4 so = x

OpenStudy (anonymous):

...What?

OpenStudy (lgbasallote):

you have [(x^3)/(x^-1)]^1/4 right? negative exponent means put the x^-1 in the numerator so [(x^3)(x^1)]^1/4 according to alws on exponents (x^3)(x^1) = x^(3+1) so that's equal to x^4. (x^4)^1/4 = x

OpenStudy (anonymous):

okay